Assume that the time a students stays in school is normally distributed with a mean of 5 hours and a standard deviation of 0.5 hours. Every day, Clark Kent stays in school for 5.5 hours. What proportion of students stays less than 5.5 hours


1
Expert's answer
2021-11-07T17:59:07-0500

Let "X=" the time a students stays in school: "X\sim N(\mu, \sigma^2)."

Given "\mu=5, \sigma=0.5."


"P(X<5.5)=P(Z<\dfrac{5.5-5}{0.5})=P(Z<1)"

"\approx0.841345"

"84.1345\%" of students stay less than 5.5 hours.